Manuscript Title: with Forced Linebreak

Ann Author^{†} and Second Author^{‡}

Authors’ institution and/or address This line break forced

Charlie Author^{§}

Second institution and/or address This line break forced

(Dated: October 11, 2000)

An article usually includes an abstract, a concise summary of the work covered at length in the main body of the article. It is used for secondary publications and for information retrieval purposes. Valid PACS numbers may be entered using the \pacs{#1} command.

PACS numbers: Valid PACS appear here

CONTENTS

I First-level heading:

The line break was forced via \\ 1 A Second-level heading:

The line break was forced via \\ …. 1

II  Displayed equations  2

A Another second-level heading . . . . . . 2

1 Third-level heading:

The line break was forced via \\ . . 2 2 Multiline equations . . . . . . . . . 2 3 Wide equations . . . . . . . . . . . 3

III Figures, Tables, and Cross-referencing 3

Acknowledgments 3

A  Appendixes  3
B  A little more on appendixes  4

1 A subsection in an appendix . . . . . . 4

References 5

I. FIRST-LEVEL HEADING:

†Also at Physics Department, XYZ University.

‡Electronic address: Second.Author@institution.edu
§URL: http://www.Second.institution.edu/~Charlie.Author

THE LINE BREAK WAS FORCED VIA \\

Here is the first sentence in Sec. I, demonstrating section cross-referencing. Here is an open-face one: 11.

A Second-level heading: The line break was forced via \\

Here is the first sentence in Sec. I A, demonstrating section cross-referencing. The environment widetext will make the text the width of the full page, as on page 3. A blank input line tells TEX that the paragraph has ended.

The width-changing commands only take effect in twocolumn style; the default is preprint style, which gives output of a constant width.

This file may be run in both preprint and twocolumn styles. Preprint format may be used for submission purposes. Twocolumn format is used to mimic final journal output.

When commands are referred to in this example file, they are always shown with their required arguments, using normal TEX format. In this format, #1, #2, etc. stand for required author-supplied arguments to commands. For example, in \section{#1} the #1 stands for the title text of the author’s section heading, and in \title{#1} the #1 stands for the title text of the paper.

Reference citations in text use the commands

\cite{#1} or \onlinecite{#1}. #1 may contain letters

and numbers. In the reference section of this paper each reference is “tagged” by the \bibitem{#1} command. #1 should be identical in both commands.

The form for citing in text is \cite{#1}, and the result is shown here [1, 2]. When needing to explicitly use the word, “Reference”, e.g., at the beginning of a sentence, use \onlinecite{#1} (Refs. 1 and 3). It is worth mentioning that REVTEX “collapses” lists of reference numbers where possible. We now cite everyone together [1–3], and once again (Refs. 1–3).

When the prb class option is used, the command \onlinecite{#1} will put the reference citations on-line: this effect appears in the preceding paragraph. Note that the location of citations must be adjusted to the reference style: the superscript references in prb style must appear after punctuation; other styles must appear before any punctuation[4]. This sample was written for the regular (non-prb) citation style, but invoking the prb option will show the results of the command \onlinecite{#1} in the preceding paragraph.

II. DISPLAYED EQUATIONS

A Another second-level heading

1 Third-level heading: The line break was forced via \\

Here is the first sentence in Sec. II A 1, demonstrating section cross-referencing. In L^{A}TEX there are many different ways to display equations, and a few preferred ways are noted below. Displayed math will center by default.

a Fourth-level heading: Single-line equations. Below we have numbered single-line equations; this is the most common type of equation in Physical Review:

χ_{+}(p) < ∼ [2|p|(|p| + p_{z})]^{−1/2} ( \begin{matrix}|p| + p_{z} \\ px + ip_{y} \end{matrix} ) , (1)
\{11234567890abc123αβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} . (2)

Note the open one in Eq. (2).

Not all numbered equations will fit within a narrow column this way. The equation number will move down automatically if it cannot fit on the same line with a one-line equation:

\{ab12345678abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} .

(3)

When the \label{#1} command is used [cf. input for Eq. (2)], the equation can be referred to in text without your knowing the equation number that TEX will assign to it. Just use \ref{#1}, where #1 is the same name that you used in the \label{#1} command.

If you have a single-line equation that you don’t want numbered, you can use the \[, \] format:

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} ….

2 Multiline equations

Multiline equations are obtained by using the

                                                  \begin{eqnarray}, \end{eqnarray} format. Use the

\nonumber command at the end of each line where you do not want a number:

M =ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}δ_{σ_{1},−σ_{2}}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{j}l_{i}ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}),  (4)

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ∑_{perm} \frac{1}{S_{12}} \frac{1}{S_{12}} ∑_{τ} c_{τ}^{f} . (5)

Note: do not use \label{#1} on a line of a multiline equation if \nonumber is also used on that line. Incorrect cross-referencing will result.

If you wish to set a multiline equation without any equation numbers, you can use the \begin{eqnarray*},

                                                  \end{eqnarray*} format:

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

× (∑_{i<j}) ( ∑_{perm} \frac{1}{S_{12}S_{23}S_{n1}} ) \frac{1}{S_{12}} .

To obtain numbers not normally produced by the automatic numbering, use the \tag{#1} command, where #1 is the desired equation number. For example, to get an equation number of (2.6′),

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} …. (2.6′)

                                                    A few notes on \eqnum{#1}. The \eqnum{#1} must

come before the \label{#1}, if any. The numbering set with \eqnum{#1} is transparent to the automatic numbering in REVTEX; therefore, you must know the number ahead of time, and must make sure that the number set with \eqnum{#1} stays in step with the automatic numbering. \eqnum{#1} works with both single-line and multiline equations. You could, if you wished, do all the numbering in a paper manually with \eqnum{#1}.

Enclosing single-line and multiline equations in

                                                  \begin{subequations} and \end{subequations} will

produce a set of equations that are “numbered” with letters, as shown in Eqs. (6a) and (6b) below:

\{abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} , (6a)

M =ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}).  (6b)

If you use a \label{#1} command right after the

                                                  \begin{subequations}, then \ref{#1} can be used to

reference all the equations in a subequations environment. For example, the equations in the preceding subequations environment were Eqs. (6).

3 Wide equations

The equation that follows is set in a wide format, i.e., it spans across the full page. The wide format is reserved for long equations that cannot be easily broken into four lines or less:

R^{(d)} = g_{σ_{2}}^{e} (\frac{[Γ^{Z}(3, 21)]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2}} + \frac{[Γ^{Z}(13, 2)]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2}} ) + x_{W} Q_{e} (\frac{[Γ^{γ}(3, 21)]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2}} + \frac{[Γ^{γ}(13, 2)]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2}} ) .  (7)

This is typed so you can see that the output is in wide format. (Since there is no input line between \end{equation} and this paragraph, there is no paragraph indent for this paragraph.) We also have

R^{(f)} = −g^{3}δ_{σ_{1},σ_{2}} (\frac{g_{σ_{2}}^{e}D_{Z}}{cos θ_{W}} − Q_{e}D_{γ} cos θ_{W} ) (\frac{[ϵ_{3}]_{σ_{1}}}{Q_{12}^{2} − M_{W}^{2} /ξ} ϵ_{1} · ϵ_{2} + \frac{[ϵ_{2}]_{σ_{1}}}{Q_{13}^{2} − M_{W}^{2} /ξ} ϵ_{1} · ϵ_{3}) .  (8)

III. FIGURES, TABLES, AND CROSS-REFERENCING

REVTEX will automatically number sections, equations, figure captions, and tables. In order to reference them in text, use the \label{#1} and \ref{#1} commands.

The \label{#1} command appears following a section heading; within an equation; or within a figure or table environment, inside of or following the caption. The \ref{#1} command appears in text where citation is to occur. We will refer to the first figure (Fig. 1) here. We can refer to the “later figure” also (Fig. 2).

References to figures: Fig. 1, Fig. 2, Fig. 3, and Fig. 4. References to tables: Table I, Table II, Table III, Table

IV, Table V, and Table VI.

Physical Review style requires that the initial citation of figures or tables be in numerical order in text, so don’t cite Fig. 3 until you’ve cited Fig. 2. See Style and Notation Guide.

ACKNOWLEDGMENTS

We wish to acknowledge the support of the author community in using REVTEX, offering suggestions and encouragement, testing new versions, ….

FIG. 1: A figure caption. The figure captions are automatically numbered.

APPENDIX A: APPENDIXES

To start the appendixes, you should use the \appendix command. This signals that all following section commands refer to appendixes instead of regular sections. Therefore, the \appendix command should be used only once—to setup the section commands to act as appendixes. Thereafter normal section commands are used. The heading for a section can be left empty. For example,

\appendix
\section{}

will produce an appendix heading that says “APPENDIX A” and

\appendix
\section{Background}

will produce an appendix heading that says “APPENDIX A: BACKGROUND” (note that the colon is set automatically).

If there is only one appendix, then the letter “A” should not appear. This is suppressed by using the star version of the appendix command (\appendix*{#1}).

FIG. 2: The “late figure.” This figure was inserted when the paper was finished. Since the figures are automatically numbered, no renumbering in text was necessary. All that needed to be done was to type the caption in the proper place and cite the figure in text.

FIG. 3: A figure caption. Use the figure* environment to get a wide figure.

TABLE III: A wide table. Two alternative occupations of special positions by KMnCL_{3} ions in the two space groups D_{4h}^{1} and D_{4h}^{1}. For a special value of the x and y parameters, a set of special positions may split into two sets of special positions of higher symmetry.

D_{4h}^{1}  D_{4h}^{5}
Ion  1st alternative  2nd alternative  lst alternative  2nd alternative
K  (2e) + (2f)  (4i)  (2c) + (2d)  (4f)
Mn  (2g)^{a}  (a) + (b) + (c) + (d)  (4e)  (2a) + (2b)
Cl  (a) + (b) + (c) + (d)  (2g)^{a}  (4e)^{a}
He  (8r)^{a}  (4j)^{a}  (4g)^{a}
Ag  (4k)^{a}  (4h)^{a}

aThe z parameter of these positions is z ∼ \frac{1}{4} .

FIG. 4: A figure caption. The labels you give tables and figures can be descriptive (as that of Fig. 1, which has a \label{autonum}) or they can reflect their numerical order, as that of this figure (\label{fig4}).

TABLE I: This is a narrow table, which occupies the entire width of a narrow column. We have used the star-form of the tabular environment to set the table on this width, and we have put \extracolsep glue between columns to provide for the necessary padding. Table captions are numbered automatically. This table illustrates left-aligned, centered, and right-aligned columns. It also shows one of two possible methods of setting footnotes within a table. In this table the footnotes are numbered and set automatically. Simply use \footnote{#1} to set a footnote mark and its text.

One^{a}  Two^{b}  Three
one  two  three
one  two  three

aNote a. bNote b.

APPENDIX B: A LITTLE MORE ON APPENDIXES

Observe that this appendix was started by using

                                                  \section{A little more on appendixes}

Note the equation number in an appendix:

E = mc^{2}.  (B1)

1 A subsection in an appendix

You can use a subsection or subsubsection in an appendix. Note the numbering: we are now in Appendix B 1.

Note the equation numbers in this appendix, produced with the subequations environment:

E = mc,  (B2a)
E = mc^{2},  (B2b)
E > ∼ mc^{3}.  (B2c)

They turn out to be Eqs. (B2a), (B2b), and (B2c).

TABLE II: This is a table of medium width. This table shows footnotes numbered manually. In this approach, \footnotemark[#1] is used to produce the footnote mark. #1 is a numeric value. Each time the same value for #1 is used, the same mark is produced in the table. After the end of the tabular environment, \footnotetext[#1]{#2} commands are used: #1 represents the same numbers used in \footnotemark[#1], and #2 represents the text of the footnote. Use these two commands to number footnotes by hand. Inspect the L^{A}TEX input for this table to see exactly how it is done.

r_{c} (Å) r_{0} (Å) κr_{0}  r_{c} (Å) r_{0} (Å) κr_{0}
Cu 0.800 14.10 2.550 Sn^{a} 0.680 1.870 3.700
Ag 0.990 15.90 2.710 Pb^{b} 0.450 1.930 3.760
Au 1.150 15.90 2.710 Ca^{c} 0.750 2.170 3.560
Mg 0.490 17.60 3.200 Sr^{d} 0.900 2.370 3.720
Zn 0.300 15.20 2.970 Li^{b} 0.380 1.730 2.830
Cd 0.530 17.10 3.160 Na^{e} 0.760 2.110 3.120
Hg 0.550 17.80 3.220 K^{e} 1.120 2.620 3.480
Al 0.230 15.80 3.240 Rb^{c} 1.330 2.800 3.590
Ga 0.310 16.70 3.330 Cs^{d} 1.420 3.030 3.740
In 0.460 18.40 3.500 Ba^{e} 0.960 2.460 3.780
Tl 0.480 18.90 3.550

aHere’s the first, from Ref. 1. bHere’s the second. cHere’s the third. dHere’s the fourth. eAnd etc.

TABLE IV: Another wide table. Numbers in columns Three–Five have been aligned by using the “d” column specifier. Non-numeric entries (those entries without a “.”) in a “d” column are aligned on the decimal point. Use the “D” specifier of the dcolumn package for more complex layouts.

One  Two  Three  Four  Five
one  two  three  four  five
He  2  2.77234  45672.  0.69
C^{a}  C^{b}  12537.64  37.66345  86.37

aSome tables require footnotes. bSome tables need more than one footnote.

TABLE V: A “late table.” This table was added after most of the paper had been completed. Since the tables are automatically numbered, no renumbering in text was necessary. This table shows the use of the the “d” column for aligning numbers on the decimal point, and the use of the “@-expression” to do finer control of alignment. This technique is useful for simple columns of numerical data, but for more complex alignments it may be necessary to use the “D” specifier of the dcolumn package or the general features of the array package.

Align by .  Multiple alignments
23.890 12  23 .890 12 ±  0 .002
12 323.  123 223  ± 344
0.834 390 12  80 .80 ± 45 .3416

TABLE VI: The Poisson ratio defined as the ratio of lateral contraction to longitudinal expansion for uniaxial stress. Experimental values are given for comparison.

σ  σ
Predicted Observed^{a}  Predicted Observed^{a}
Cu  0.48  0.36  Al  0.47  0.33
Ag  0.48  0.37  Tl  0.47  0.35
Au  0.48  0.36  Sn  0.46  0.33
Mg  0.47  0.35  Pb  0.46  0.40–0.45
Zn  0.47  0.25  Pb  0.49  0.43
K  0.49  0.44

REFERENCES

[1] A. Smith and B. Doe, J. Chem. Phys. 76, 4056 (1982).

[2] C. Jones, J. Chem. Phys. 68, 5298 (1978).

[3] C. Jones and A. Smith, J. Chem. Phys. 72, 5168 (1980).

[4] Authors are encouraged to use BibTEX and prsty.bst to create their reference list in proper APS style. Instructions can be requested by e-mail (mailto:

                                                     mis@aps.org).